# DeepShade: Enable Shade Simulation by Text-conditioned Image Generation

### v1

@@ Abstract
Heatwaves pose a significant threat to public health, especially as global warming intensifies. However, current routing systems (e.g., online maps) fail to incorporate shade information due to the difficulty of estimating shades directly from noisy satellite imagery and the limited availability of training data for generative models. In this paper, we address these challenges through two main contributions. First, we build an extensive dataset covering diverse longitude-latitude regions, varying levels of building density, and different urban layouts. Leveraging Blender-based 3D simulations alongside building outlines, we capture building shadows under various solar zenith angles throughout the year and at different times of day. These simulated shadows are aligned with satellite images, providing a rich resource for learning shade patterns. Second, we propose the DeepShade, a diffusion-based model designed to learn and synthesize shade variations over time. It emphasizes the nuance of edge features by jointly considering RGB with the Canny edge layer, and incorporates contrastive learning to capture the temporal change rules of shade. Then, by conditioning on textual descriptions of known conditions (e.g., time of day, solar angles), our framework provides improved performance in generating shade images. We demonstrate the utility of our approach by using our shade predictions to calculate shade ratios for real-world route planning in Tempe, Arizona. We believe this work will benefit society by providing a reference for urban planning in extreme heat weather and its potential practical applications in the environment.
@@ 1 Introduction
Extreme weather is causing an increasing number of deaths worldwide, with heatwaves being a major contributing factor. According to a report Ritchie (2024), the frequency and intensity of extreme heat events have surged over the past two decades, with more than 178,700 deaths occurring annually (average from 2000 to 2019) as a direct result of high temperatures Shanshan Li (2021). Research from the World Health Organization highlights that extreme heat is now one of the leading causes of weather-related deaths NIEHS (2022), disproportionately affecting vulnerable populations such as the elderly and those working or staying in urban areas with limited access to cooling infrastructure. This trend underscores the urgent need for adaptive measures, including heat-resilient urban design and shade-aware route-planning methods Ma (2018), to mitigate the public health impact of rising temperatures globally.
@@
Since shade acts as a natural shelter that reduces direct exposure to solar radiation, understanding how shade changes in real-time is crucial for preparing outdoor activities and aiding urban planning in establishing artificial shelters in areas lacking natural shade. As introduced in Da et al. (2024a), by identifying shaded areas and integrating this information into route-planning systems, individuals can make safer travel choices, reducing their risk of heat-related illnesses. However, the existing research faces significant limitations: First, shade analysis using urban simulations is only on static maps, which lacks good generalizability. Second, most of the methods are localized, relying on resource-intensive LiDAR data; they also lack scalability across different or more extensive areas. Third, they are unable to capture the real-time shade dynamics, limiting their utility in intelligent routing. These limitations impede the development of accurate shade modeling and timely planning, affecting the practical impact of existing studies.
@@
To tackle the above challenges, a more adaptive and scalable approach to dynamically model shade variations is required. Given the success of Generative AI Da et al. (2024b) (particularly diffusion models), in capturing spatial-temporal patterns in urban scene synthesis and environment simulation Da et al. (2025b), they present a promising direction for overcoming these challenges by generating shades for satellite images. However, to effectively leverage this method, a well-structured dataset, that accurately aligns ground-truth shade-variance with its satellite-image level geographic information, is necessary. Such a dataset is left blank in the current research domain.
@@
Figure 1: The overview of the pipeline, note that ‘Addr.’ is an abbreviation of Address, and ‘Roof*’ includes: roof shape and roof height, the building shape is approximated by the bird-eye view of the roof shape.
@@
In this paper, we first developed a rigorous and systematic pipeline to construct a comprehensive dataset for training shade-generation models. The dataset is designed to encompass three critical dimensions while adhering to a unified standard. The three dimensions include: (1) Geographical Diversity, covering a wide range of continents with varying latitudinal and longitudinal distributions; (2) Urban Layout Variability, capturing diverse urban configurations such as dense high-rise building places and sparse flat areas; and (3) Traffic Rule Variationaccounting for differences in driving orientations, including left-hand and right-hand traffic systems. The unified standard is that the cities must suffer from significant heat events as defined by the World Meteorological Organization (WMO), which guarantees the dataset focuses on regions where high temperatures and their associated impacts on shade dynamics are most pronounced. We secondly, designed a novel contrastive learning-based diffusion model approach, with a fine-granularity edge conditional module, that learns the shade variations based on the skeleton representations of corresponding satellite images. The model effectively maps shade dynamics to text prompts containing temporal and geographic information, allowing it to generate realistic shade predictions for arbitrary times of day or specific solar angles. This solution is unique for its ability to train on a limited satellite image dataset while generalizing well to unseen buildings, provided a satellite image is available. This relaxed the requirements for detailed building height to conduct simulation and ensures adaptability and robustness across diverse urban settings.
@@ 3 Dataset Creation
In order to develop generalizable models, we construct a formal pipeline that creates a comprehensive dataset covering three dimensions: Geographical Diversity, Urban Layout Variability, and Traffic Rule Variation. This dataset aligns satellite images with Open Street Map (OSM) Map (2017) building information. The pipeline involves four major steps as shown in Figure 1, which illustrates the overall framework from input to the collected outcome.
First, Metadata Retrieval obtains all necessary metadata used for the 3-D simulation. To align with the research community and utilize the large open-source data, we adopt OSM data as the metadata for shade simulation. Moreover, based on the longitude and latitude information, we can extract the building’s geographical information including “property type, address, height, levels, shape, and geometry set of locations”, this typically provides information such as a list of points that forms a polygon of building’s 3D skeleton.
Second, Simulation Setups. With the necessary data collected, we performed the simulation to capture the shade changes for our dataset creation. In addition, we use Blender Gschwandtner et al. (2011) for large-scale city-wide shade simulation. Based on scientific rules, we set up a controller as the sun’s movement engine, which can generate the sun’s trajectory following the sun’s solar declination or angle on any day of the year, and at any time in a day. Then, we rendered the area of interest (AOI) within the simulation platform and adjust the scale to ensure it aligns with the prevalent maps, here, we align images with Google map tile level 13.
@@
The shaded snapshot $x^{shade}$ is obtained with parallel sunshine casting on the scene, capturing shades, side projections, and full structural visibility of buildings. To extract a structure-only reference, we generate a skeleton snapshot $x^{sk}$ by rendering the identical scene with the sun being invisible. Additionally, a satellite image $x^{sat}$ of the corresponding area is aligned the simulated environment with real-world conditions. Finally, we extract text descriptions $T$ encoding temporal information, such as the sun angle and time of day, which provide an additional modality for shade generation control. We formulate the text description as follow:
@@
Prompt 1: Solar declination: -20.7
Prompt 2: Angle: 45
@@
Last, we have the Post process step. To benefit the deep learning models, we need a ground-truth shade for measuring the model’s output, and provide feedback by loss functions. However, current data from step 3, are not the ultimate shades as they include side shadows and noises. Therefore, to obtain the actual ground truth shade, we propose to take another snapshot of the scene without sunshine, named as ‘skeleton image’ $x^{sk}$, which only provides a pure building structure, then the pure shade can be obtained through formula:
@@
where $x^{gt}$ represents the extracted ground truth shade, $x^{shade}$ is the shaded snapshot taken under sunlight, capturing the shades, sides, and full building observation, and $x^{sk}$ is the skeleton snapshot taken without sunlight, providing only the pure building structure. The term $\mathbb{I}(x^{shade}\leq\alpha)$ is an indicator function that removes low-intensity grey values (noise thresholding), where $\alpha$ is a predefined thresholdThis equation effectively removes structural elements and noise, isolating the actual shaded areas to provide precise ground truth data for machine learning models.
The dataset construction process has improved alignment between simulated and real-world data while enabling controllable shade generation via images, textual descriptions, or a joint approach. This dataset has been split into train, test sets by 70%, 30%, which can help the research domain with a well-defined benchmark (use could hold another validation set by changing the index if necessary).
@@ 4 A Text Conditioned Shaded Image Generation Model
Now given a well-constructed shade dataset $D=\{x^{shade},x^{sk},x^{sat},T\}$, where $x^{shade}$ is the shaded snapshot capturing shadowed areas under sunlight, $x^{sk}$ is the skeleton snapshot, $x^{sat}$ is the real-world satellite image after alignment, and $T$ is the textual description encoding solar angle $\theta_{sun}$ and timestamp $t_{day}$. To benefit from that, this paper also introduces a novel method that generates accurate shade images following the text prompt and base image (which provides a base map for generation). Given the objective is to generate shade images using a base map and text conditions, we adopt a diffusion-based framework, ControlNet Zhang et al. (2023), as our backbone since the diffusion models are excellent in image synthesis and editing tasks Dhariwal and Nichol (2021). As shown in Figure 2, our proposed method focuses on edge information incorporation and contrastive learning module. We will introduce the details after a reflection on the Controlnet Model in our setting.
@@
The vanilla ControlNet framework could functionally generate shade images but the quality of generation is limited, mainly because of two reasons:
@@ 4.1 Edge‐Enhanced Conditional Generation
To address the first challenge that the vanilla model fails to capture subtle shade outline features and often generates shadows with crooked boundaries even for buildings with straight outlines, we propose an edge‐enhanced conditional generation method. The approach incorporates edge information as an additional visual condition alongside the base map, as shown in the preparation part of Figure 2.
Specifically, our base map $x^{sk}\in\mathbb{R}^{H\times W\times 3}$ is a three‐channel mask of building footprints. We extract structural boundaries via Canny edge detection to obtain a single‐channel edge map $x^{edge}\in\mathbb{R}^{H\times W\times 1}$. These are concatenated along the channel dimension to form a four‐channel conditioning tensor:
$x^{cond}_{ij}=\bigl{[}x^{sk}_{ij,R},\,x^{sk}_{ij,G},\,x^{sk}_{ij,B},\,x^{edge}_{ij}\bigr{]}\;\in\;\mathbb{R}^{H\times W\times 4}.$ (4)
This four‐channel tensor is supplied to the ControlNet U-Net as its concatenated input, together with the text prompt $T$. By providing both the global building layout and its precise edge contours, the generator as shown below:
@@ 4.2 Contrastive Based Shade Generation
Baselines on Scenario1 Beijing (CHN) Phoenix (USA) Sau Paulo (BRA) Madrid (ESP) Cairo (EGY) Mumbai (IND)
@@ 4.2.1 Contrastive Buffer Pairs Creation
To effectively train a model capable of generating shades following the temporal rules (the generated shades between nearby time steps should be more similar than those with larger temporal gaps), we design a contrastive learning framework. This section introduces how we construct a dataset of contrastive pairs: positive and negative pairs, based on their temporal and spatial relationships.
@@ 4.2.2 Contrastive Learning for Shade Generation
During training, for each image pair $(x_{i},x_{j})$, feature embeddings $h_{i}$ and $h_{j}$ are obtained from a pre-trained embedding network using U-net  [^2] The reason for pre-training is introduced in discussion part,
These embeddings are used to compute a similarity matrix $S\in\mathbb{R}^{N\times N}$, where $N=2\times B$ is the total number of embeddings in the batch (each pair contributing two embeddings), and each element $S_{uv}$ represents the cosine similarity between embeddings as below:
@@ 5 Experiment
In this section, we conduct experiments to verify the effectiveness of our proposed method. We have designed three sets of experiments, the first is to verify the model’s performance in dense building cities, such as Beijing, Phoenix downtown, and Sau Paulo, The second is to show the performance in the relatively sparse environment such as Tempe, Brasilia, etc. Please note that these cities’ selection is based on their representation in each continent and also because they suffer from extreme heatwaves in summer as history records. The third, we perform an ablation study to understand the contribution of different components in our methods.
@@ 5.1 Metrics
Table 2: Comparison of different models across various metrics: SSIM, MSE, mIoU, LPIPS, and B-IoU. In this experiment, each of the trained models is fed with a bird-eye view satellite skeleton image and a text prompt describing the time and solar angle, and our method consistently performs better than these baseline methods.
@@ 5.2 Experiment Analysis
In this part, we focus on the Table. 1. The upper part is the dense areas in the world selected dataset for testing, and the lower part is the sparse areas. We can observe that the DeepShade outperforms most of the baseline methods in the various metrics, no matter of the dense or relatively sparse scenario land covers. It demonstrates the vivid simulation of the conditioned area using the ControlNet-based method by training on one dataset, with this good transferability, it better indicates that this work has great potential for real-world applications.
Ablation Study: We also included an ablation study in the paper, as in Table 2, it is the training conducted in the Tempe dataset, and the test is the other 30% from the original dataset given by the default split. This result reveals the importance of each component in our method. We can see that the edge condition (our method without contrastive learning) suffers the most performance drop in comparison to the DeepShape full model structure, and if we also remove the edge conditions, the performance further decreases, the backbone model is the stable diffusion model (all of the above models are trained 5 times with mean and std reported).

### v2

@@ Abstract
Heatwaves pose a significant threat to public health, especially as global warming intensifies. However, current routing systems (e.g., online maps) fail to incorporate shade information due to the difficulty of estimating shades directly from noisy satellite imagery and the limited availability of training data for generative models. In this paper, we address these challenges through two main contributions. First, we build an extensive dataset covering diverse longitude-latitude regions, varying levels of building density, and different urban layouts. Leveraging Blender-based 3D simulations alongside building outlines, we capture building shadows under various solar zenith angles throughout the year and at different times of day. These simulated shadows are aligned with satellite images in terms of the areas, providing a rich resource for learning shade patterns. Second, we propose the DeepShade, a diffusion-based model designed to learn and synthesize shade variations over time. It emphasizes the nuance of edge features by jointly considering RGB with the Canny edge layer, and incorporates contrastive learning to capture the temporal change rules of shade. Then, by conditioning on textual descriptions of known conditions (e.g., time of day, solar angles), our framework provides improved performance in generating shade images. We demonstrate the utility of our approach by using our shade predictions to calculate shade ratios for real-world route planning in Tempe, Arizona. We hope this work could provide a reference for urban planning in extreme heat weather and reveal its potential practical applications in the environment.
@@ 1 Introduction
Extreme weather is causing an increasing number of deaths worldwide, with heatwaves being a major contributing factor. According to a report Ritchie (2024), the frequency and intensity of extreme heat events have surged over the past two decades, with more than 178,700 deaths occurring annually (average from 2000 to 2019) as a direct result of high temperatures Shanshan Li (2021). Research from the World Health Organization highlights that extreme heat is now one of the leading causes of weather-related deaths NIEHS (2022), disproportionately affecting vulnerable populations such as the elderly and those working or staying in outdoor areas with limited access to cooling infrastructure. This trend underscores the urgent need for adaptive measures, including heat-resilient urban design and shade-aware route-planning methods Ma (2018), to mitigate the public health impact of rising temperatures globally.
@@
Since shade acts as a natural shelter that reduces direct exposure to solar radiation, understanding how shade changes in real-time is crucial for preparing outdoor activities and aiding urban planning in establishing artificial shelters in areas lacking natural shade. As introduced in Da et al. (2024a), by identifying shaded areas and integrating this information into route-planning systems, individuals can make more comfortable travel choices, reducing their risk of heat-related illnesses. However, the existing research faces significant limitations: First, shade analysis using urban simulations is only on static maps, which lacks good generalizability. Second, most of the methods are localized, relying on resource-intensive LiDAR data; they also lack scalability across different or more extensive areas. Third, they are unable to capture the real-time shade dynamics, limiting their utility in intelligent routing. These limitations impede the development of accurate shade modeling and timely planning, affecting the practical impact of existing studies.
@@
To tackle the above challenges, a more adaptive and scalable approach to dynamically model shade variations is required. Given the success of Generative AI Da et al. (2024b) (particularly diffusion models), in capturing spatial-temporal patterns in urban scene synthesis Tang et al. (2024) and environment simulation Da et al. (2025b), they present a promising direction for overcoming these challenges by generating shades for satellite images. However, to effectively leverage this method, a well-structured dataset, that accurately aligns ground-truth shade-variance with its satellite-image level geographic information, is necessary. Such a dataset is left blank in the current research domain.
@@
Figure 1: The overview of the pipeline, note that ‘Addr.’ is an abbreviation of Address, and ‘Roof*’ includes: roof shape and roof height, the building shape is approximated by the bird’s-eye view of the roof shape.
@@
In this paper, we first developed a rigorous and systematic pipeline to construct a comprehensive dataset for training shade-generation models. The dataset is designed to encompass three critical dimensions while adhering to a unified standard. The three dimensions include: (1) Geographical Diversity, covering a wide range of continents with varying latitudinal and longitudinal distributions; (2) Urban Layout Variability, capturing diverse urban configurations such as dense high-rise building places and sparse flat areas; and (3) Traffic Rule Variation, accounting for differences in driving orientations, including left-hand and right-hand traffic systems. The unified standard is that the cities must suffer from significant heat events as defined by the World Meteorological Organization (WMO), which guarantees the dataset focuses on regions with high temperatures and their associated impacts on shade dynamics are most pronounced. We secondly, designed a novel contrastive learning-based diffusion model approach, with a fine-granularity edge conditional module, that learns the shade variations based on the skeleton representations of corresponding satellite images. The model effectively maps shade dynamics to text prompts containing temporal and geographic information, allowing it to generate realistic shade predictions for arbitrary times of day or specific solar angles. This solution is unique for its ability to train on a limited satellite image dataset while generalizing to unseen buildings, provided a satellite image is available. This relaxed the requirements for detailed building height to conduct simulation and ensures adaptability and robustness across diverse urban settings.
@@ 3 Dataset Creation
In order to develop generalizable models, we construct a formal pipeline that creates a comprehensive dataset covering three dimensions: Geographical Diversity, Urban Layout Variability, and Traffic Rule Variation. This dataset aligns satellite images with Open Street Map (OSM) Ramm et al. (2025) building information. The pipeline involves four major steps as shown in Figure 1, which illustrates the overall framework from input to the collected outcome.
First, Metadata Retrieval obtains all necessary metadata used for the 3-D simulation. To align with the research community and utilize the large open-source data, we adopt OSM data as the metadata for shade simulation. Moreover, based on the longitude and latitude information, we can extract the building’s geographical information, including “property type, address, height, levels, shape, and geometry set of locations”, this typically provides information such as a list of points that form a polygon of building’s 3D skeleton.
Second, Simulation Setups. With the necessary data collected, we performed the simulation to capture the shade changes for our dataset creation. In addition, we use Blender Hess (2013) for large-scale city-wide shade simulation. Based on scientific rules, we set up a controller as the sun’s movement engine, which can generate the sun’s trajectory following the sun’s solar declination or angle on any day of the year, and at any time of the day. Then, we rendered the area of interest (AOI) within the simulation platform and adjusted the scale to ensure it aligns with the prevalent maps. Here, we align images with Google map tile level 13.
@@
The shaded snapshot $x^{shade}$ is obtained with parallel sunshine casting on the scene, capturing shades, side projections, and full structural visibility of buildings. To extract a structure-only reference, we generate a skeleton snapshot $x^{sk}$ by rendering the identical scene with the sun being invisible. Additionally, a satellite image $x^{sat}$ of the corresponding area is aligned the simulated environment with real-world conditions. Finally, we extract text descriptions $T$ encoding temporal information, such as the sun angle and time of day, which provide an additional modality for shade generation control. We formulate the text description as follows:
@@
Prompt 1: Solar declination: -20.7°
Prompt 2: Angle: 45°
@@
Last, we have the Post process step. To benefit the deep learning models, we need a ground-truth dataset for measuring the model’s output and providing feedback through loss functions. However, current data from step 3 is not the ultimate shade as it includes side shadows and noises. Therefore, to obtain the actual ground truth shade, we propose to take another snapshot of the scene without sunshine, named as ‘skeleton image’ $x^{sk}$, which only provides a pure building structure, then the pure shade can be obtained through the formula:
@@
where $x^{gt}$ represents the extracted ground truth shade, $x^{shade}$ is the shaded snapshot taken under sunlight, capturing the shades, sides, and full building observation, and $x^{sk}$ is the skeleton snapshot taken without sunlight, providing only the pure building structure. The term $\mathbb{I}(x^{shade}\leq\alpha)$ is an indicator function that removes low-intensity grey values (noise thresholding), where $\alpha$ is a predefined threshold. This equation effectively removes structural elements and noise, isolating the actual shaded areas to provide precise ground truth data for machine learning models.
The dataset construction process has improved alignment between simulated and real-world data while enabling controllable shade generation via images, textual descriptions, or a joint approach. This dataset has been split into train, test sets by 70%, 30%, which can help the research community with a well-defined benchmark (researcher could hold another validation set by changing the index if necessary).
@@ 4 A Text Conditioned Shaded Image Generation Model
Now given a well-constructed shade dataset $D=\{x^{shade},x^{sk},x^{sat},T\}$, where $x^{shade}$ is the shaded snapshot capturing shadowed areas under sunlight, $x^{sk}$ is the skeleton snapshot, $x^{sat}$ is the real-world satellite image after alignment, and $T$ is the textual description encoding solar angle $\theta_{sun}$ and timestamp $t_{day}$. To benefit from that, this paper also explores methods that generate accurate shade images following the text prompt and base image (which provides a base map for generation). Given the objective is to generate shade images using a base map and text conditions, we adopt a diffusion-based framework, ControlNet Zhang et al. (2023), as our backbone since the diffusion models are effective in image synthesis and editing tasks Dhariwal and Nichol (2021). As shown in Figure 2, our proposed method focuses on edge information incorporation and contrastive learning module. We will introduce the details after a reflection on the Controlnet Model in our setting.
@@
The vanilla ControlNet framework could functionally generate shade images, but the quality of generation is limited, mainly because of two reasons:
@@ 4.1 Edge‐Enhanced Conditional Generation
To address the first challenge that the vanilla model fails to capture subtle shade outline features and often generates shadows with crooked boundaries, even for buildings with straight outlines, we propose an edge‐enhanced conditional generation method. The approach incorporates edge information as an additional visual condition alongside the base map, as shown in the preparation part of the Figure 2.
Specifically, our base map $x^{sk}\in\mathbb{R}^{H\times W\times 3}$ is a three‐channel mask of building footprints. We extract structural boundaries via Canny edge detection to obtain a single‐channel edge map $x^{edge}\in\mathbb{R}^{H\times W\times 1}$. These are concatenated along the channel dimension to form a four‐channel conditioning tensor (R, G, B and edge):
$x^{cond}=\bigl{[}x^{sk}_{R},\,x^{sk}_{G},\,x^{sk}_{B},\,x^{edge}\bigr{]}\;\in\;\mathbb{R}^{H\times W\times 4}.$ (4)
This four‐channel tensor is supplied to the ControlNet’s U-Net as its concatenated input, together with the text prompt $T$. By providing both the global building layout and its precise edge contours, the generator is shown below:
@@ 4.2 Contrastive Based Shade Generation
Baselines on Scenario1 Beijing (CHN) Phoenix (USA) São Paulo (BRA) Madrid (ESP) Cairo (EGY) Mumbai (IND)
@@ 4.2.1 Contrastive Buffer Pairs Creation
To effectively train a model capable of generating shades following the temporal rules (the generated shades between nearby time steps should be more similar than those with larger temporal gaps), we design a contrastive learning framework. This section introduces how we construct a training set of contrastive pairs: positive and negative pairs, based on their temporal and spatial relationships.
@@ 4.2.2 Contrastive Learning for Shade Generation
During training, for each image pair $(x_{i},x_{j})$, feature embeddings $h_{i}$ and $h_{j}$ are obtained from a pre-trained embedding network using U-net, these embeddings are used to compute a similarity matrix $S\in\mathbb{R}^{N\times N}$, where $N=2\times B$ is the total number of embeddings in the batch (each pair contributing two embeddings), and each element $S_{uv}$ represents the cosine similarity between embeddings as below:
@@ 5 Experiment
In this section, we conduct experiments to verify the effectiveness of our proposed method. We have designed three sets of experiments, the first is to verify the model’s performance in dense building cities, such as Beijing, Phoenix downtown, and São Paulo, The second is to show the performance in the relatively sparse environment, such as Tempe, Brasilia, etc. Please note that these cities’ selection is based on their representation in each continent, and also because they suffer from extreme heatwaves in summer, as history records. The third, we perform an ablation study to understand the contribution of different components in our methods.
@@ 5.1 Metrics
Table 2: Comparison of different models across various metrics: SSIM, MSE, mIoU, LPIPS, and B-IoU. In this experiment, each of the trained models is fed with a bird’s-eye view satellite skeleton image and a text prompt describing the time and solar angle, and our method consistently performs better than these baseline methods.
@@ 5.2 Experiment Analysis
In this part, we focus on the Table. 1. The upper part is the dense areas in the world selected dataset for testing, and the lower part is the sparse areas. We can observe that  our model outperforms most of the baseline methods in the various metrics (with a total of 50 epochs of training), regardless of the dense or relatively sparse scenario land covers. It demonstrates the vivid simulation of the conditioned area using the ControlNet-based method by training on one dataset, with good transferability, it better indicates that this work has great potential for real-world applications.
C. Ablation Study.
We also included an ablation study in the paper, as in Table 2, it is the training conducted in the Tempe dataset, and the test is the other 30% from the original dataset given by the default split. This result reveals the importance of each component in our method. We can see that the edge condition (our method without contrastive learning) suffers the most performance drop in comparison to the DeepShape full model structure, and if we also remove the edge conditions, the performance further decreases, the backbone model is the stable diffusion model (all of the above models are trained 5 times with mean and std reported).